OpenStudy (anonymous):

graph.. x = 3 sin^3t y = 3 cos^3t

OpenStudy (mathmale):

Note that there is ambiguity in your post. Do you mean "the cubes of cos t and sin t"? or do you mean cos ^ (3t) and sin ^ (3t)? Please clear this up to the best of your ability. You could make a table of values and let t be your independent variable. Let t=0, . Calculate the coordinates (x,y) for each t value to form a single point. Plot these points.

OpenStudy (anonymous):

it is the cubes of of cos t and sin t \[3 \sin ^{3} t\]

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Nnesha

OpenStudy (mathmale):

OK. Then\[x=3\sin^3t,y=3\cos^3t\]

OpenStudy (mathmale):

What are x and y when t=0? What would x and y be were t=pi/6?

OpenStudy (anonymous):

hm.. how do i find that again?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@dan815

OpenStudy (mathmale):

\[x=3\sin^3t,y=3\cos^3t\] Evaluate these formulas for x and y at t=0, t=pi/6, y=pi/3, and so on. Example: at t=0, x=3*(sin 0)^3= 3*(0)^3 = 0.